# Think Twice Before You Judge: Mixture of Dual Reasoning Experts for Multimodal Sarcasm Detection

### v1

@@ 3.5. Model Training and Prediction
The output $H_{\hat{T}}^{L}$ from the final encoder layer $L$ is fed to the LLM decoder, which generates the label from the label set $\{yes,no\}$. The LLM and the image and text projection layers are trained to minimize the cross-entropy loss between the ground truth label and the generated label. The CLIP visual and text encoders are only used as feature extractors and are frozen during training.
@@ 4.1. Datasets and Experimental Settings
We evaluate our proposed model MiDRE on MMSD2.0 (Qin et al., 2023) dataset. We also report the performance on MMSD(Cai et al., 2019), which is a biased dataset in Table 3. The dataset statistics are reported in Table 1. We use flan-t5-large 111google/flan-t5-large as our backbone LLM and for generating rationales, we use the LVLM Qwen2-VL-7B-Instruct [^2] Qwen/Qwen2-VL-7B-Instruct, due to its better zero-shot performance on our task over other LVLMs as validated in Figure 6. For image and text feature extractors, we use clip-vit-large-patch14333openai/clip-vit-large-patch14. We train our model for 20 epochs with Adam optimizer, maintaining a batch size of 12 and a learning rate of 5e-5. We pick the model with the best validation accuracy for testing. All experiments are carried out on 1xNvidia A100 80GB GPU.
@@
Negative 1174/10230 1451/1368 1450/1372
Figure 4. Unique test samples correctly classified by MiDRE(w/o IR) vs MiDRE(w/o ER).
Table 2. Performance comparison of different models on the MMSD2.0 dataset. * denotes that we reimplemnt the models. indicates our method significantly outperforms baselines with $p<0.001$.
Model Acc. P R F1 Text-Only TextCNN (Kim, 2014) 71.61 64.62 75.22 69.52 BiLSTM (Graves and Schmidhuber, 2005) 72.48 68.02 68.08 68.05 SMSD (Xiong et al., 2019) 73.56 68.45 71.55 69.97 RoBERTa (Liu et al., 2019) 79.66 76.74 75.70 76.21 ChatGLM2-6B (Du et al., 2021) 78.41 78.15 78.65 78.23 LLaMA2-7B (Touvron et al., 2023) 82.52 82.15 82.46 82.27 Image-Only ResNet (He et al., 2015) 65.50 61.17 54.39 57.58 ViT (Dosovitskiy et al., 2020) 72.02 65.26 74.83 69.72 Multimodal HFM (Cai et al., 2019) 70.57 64.84 69.05 66.88 Att-BERT (Pan et al., 2020) 80.03 76.28 77.82 77.04 CMGCN (Liang et al., 2022) 79.83 75.82 78.01 76.90 HKE (Liu et al., 2022b) 76.50 73.48 71.07 72.25 DynRT-Net (Tian et al., 2023) 71.40 71.80 72.17 71.34 Multi-view CLIP* (Qin et al., 2023) 85.14 80.18 88.21 84.00 RAG-LLaVA* (Tang et al., 2024) 85.11 80.02 87.34 83.51 MiDRE (ours) 86.18 86.62 89.00 87.79
@@ 4.2. Baselines
We consider three groups of baselines: image-modality, text-modality and multimodal methods. For image-modality, we follow (Cai et al., 2019) and use the representations from the pooling layer of ResNet (He et al., 2015) and embedding of the [CLS] token from ViT (Dosovitskiy et al., 2020). For text-modality methods, we consider TextCNN (Kim, 2014), BiLSTM (Graves and Schmidhuber, 2005), SMSD (Xiong et al., 2019) which uses a low rank bilinear pooling using a self-matching network, RoBERTa(Liu et al., 2019) and more recent LLMs like LLaMA2-7B (Touvron et al., 2023) and ChatGLM2-6B (Du et al., 2021). For the multimodal methods, we consider HFM (Cai et al., 2019), D&R Net (Xu et al., 2020), Att-BERT (Pan et al., 2020), InCrossMGs (Liang et al., 2021), CMGCN (Liang et al., 2022), HKE (Liu et al., 2022a), DynRT-Net (Tian et al., 2023), Multi-view CLIP (Qin et al., 2023) and RAG-LLaVA (Tang et al., 2024). We do not show the results of InCrossMGs and D&R Net on MMSD2.0 due to the unavailability of their codes.
Figure 5. Examples demonstrating the importance of external rationales for sarcasm identification. ’Baseline’ indicates Multi-view CLIP (the best performing baseline for MMSD2.0), ER(IR) Gate Values represent the importance given to external (internal) expert module by the gating network in MiDRE.
@@ 4.3. Main Results
Following (Cai et al., 2019), we use Accuracy (Acc.), macro-average precision (P), macro-average recall (R), and macro-average F1 (F1) to report the performance of our model in Table 5. Our main observations are follows: (1) Our method MiDRE outperforms all SOTA baselines in MMSD2.0: On MMSD2.0 dataset, MiDRE outperforms all baseline across unimodal and multimodal settings. Specifically it ourpeforms the best SOTA multimodal baseline Multi-view CLIP by good margins across metrics (Acc: $\uparrow 1.04\%$, P: $\uparrow 6.44\%$, R: $\uparrow 0.79\%$, F1: $\uparrow 3.79\%$ ). This proves the efficacy of using structured rationales and having dual reasoning paths to identify sarcasm.
(2) MiDRE outperforms all multimodal baselines that use external knowledge: We notice that our method significantly outperform baselines like HFM (Cai et al., 2019) (Acc: $\uparrow 15.61\%$ in MMSD2.0) which uses image attributes generated using ResNet, CMGCN (Liang et al., 2022) (Acc: $\uparrow 6.35\%$ in MMSD2.0) which utilizes object-attribute pairs from a detectron module and HKE (Liu et al., 2022a) (Acc: $\uparrow 9.68\%$ in MMSD2.0) which incorporates image caption, image attributes and adjective-noun pairs from image. This highlights the importance of leveraging structured, verbose and sarcasm-oriented knowledge.
@@ 4.4. Results on Biased MMSD Dataset
MMSD is a text-biased dataset, as demonstrated by Qin et al. (2023). This bias arises from the presence of strong textual cues, such as hashtags and emojis that predominantly occur in one class. Since text-only models do not process visual information, they can directly exploit these cues without interference from the image modality, often outperforming multimodal models on this dataset. For completeness, we evaluate MiDRE on MMSD. As shown in Table 3, the text-only LLM ChatGLM2-6B even outperforms existing multimodal models. Notably, traditional language models like RoBERTa perform on par with larger LLMs such as LLaMA2-7B and ChatGLM2-6B, further emphasizing the dataset’s inherent bias. Despite this, MiDRE achieves consistently better performance than all prior multimodal approaches by significant margins, demonstrating its robust reasoning ability. Moreover, the small performance gap between MiDRE and ChatGLM2-6B suggests that MiDRE effectively leverages strong textual signals while also integrating complementary visual and external knowledge.
@@ 5. Comparison with LVLMs
We compare the performance of state-of-the-art large vision-language models (LVLMs) under a zero-shot prompting setup for MMSD2.0, as reported in Figure 6. Despite advanced zero-shot capabilities and contextual knowledge, these models perform poorly on our task. Since fine-tuning of such large models is expensive, we press upon the need of developing smaller models like MiDRE that are explicitly designed for nuanced reasoning tasks such as multimodal sarcasm detection. At the same time, it reveals how the capabilities of such smaller models can be significantly amplified by strategically leveraging contextual understanding from large vision-language models.
@@ 6. Ablation
To further prove the efficacy of our approach, we answer the following research questions: (1) What is the impact of having only one expert network? (2) What advantage does the gating network provide? (3) Do we need rationales from image, text and multimodal views?
@@ 6.1. Single Expert Network
To better understand the contributions of the internal (IR) and external (ER) reasoning experts, we perform an ablation study on MiDRE by selectively removing each component (Table 4, rows 3 & 4). Removing the ER module (w/o ER) leads to a significant drop in accuracy (↓4.72% on MMSD, ↓3.91% on MMSD2.0), highlighting the importance of external contextual rationales that are not directly inferable from the input. Conversely, removing IR (w/o IR) also reduces performance (↓2.75% on MMSD, ↓2.12% on MMSD2.0), showing that ER alone struggles to compensate for noisy or incomplete rationales.
Interestingly, from Figure 4, we can observe that w/o ER correctly classifies 111 unique samples and w/o IR classifies 154 different ones, suggesting that both experts offer complementary and independently valuable reasoning capabilities. Moreover, the larger performance drop without ER indicates its dominant role, but IR remains essential for grounding predictions in the raw image-text content, enabling more reliable sarcasm detection through their combined use.
Table 3. Performance comparison of different models on the MMSD dataset. * denotes that we reimplement the models.
Model Acc. P R F1 Text-Only TextCNN (Kim, 2014) 80.03 74.29 76.39 75.32 BiLSTM (Graves and Schmidhuber, 2005) 81.90 76.66 78.42 77.53 SMSD (Xiong et al., 2019) 80.90 76.46 75.18 75.82 RoBERTa (Liu et al., 2019) 93.97 90.39 94.59 92.45 ChatGLM2-6B (Du et al., 2021) 94.02 93.46 94.14 93.76 LLaMA2-7B (Touvron et al., 2023) 93.97 93.42 94.09 93.72 Image-Only ResNet (He et al., 2015) 64.76 54.41 70.80 61.53 ViT (Dosovitskiy et al., 2020) 67.83 57.93 70.07 63.40 Multimodal HFM (Cai et al., 2019) 83.44 76.57 84.15 80.18 D&R Net (Xu et al., 2020) 84.02 77.97 83.42 80.60 Att-BERT (Pan et al., 2020) 86.05 80.87 85.08 82.92 InCrossMGs (Liang et al., 2021) 86.10 81.38 84.36 82.84 CMGCN (Liang et al., 2022) 86.54 - - 82.73 HKE (Liu et al., 2022b) 87.36 81.84 86.48 84.09 DynRT-Net (Tian et al., 2023) 93.59 93.06 93.60 93.31 Multi-view CLIP* (Qin et al., 2023) 88.33 82.66 88.65 88.55 RAG-LLaVA* (Qin et al., 2023) 89.97 89.26 89.58 89.42 MiDRE (ours) 93.81 93.11 93.84 93.47
Figure 6. Zero-shot Performance of LVLMs on MMSD2.0 test samples.
Figure 7. Examples demonstrating how noisy rationales can provide contextual cues. ’Baseline’ indicates Multi-view CLIP (the best performing baseline for MMSD2.0), ER(IR) Gate Values represent the importance given to external (internal) expert module by the gating network in MiDRE.
@@ 6.2. Impact of Gating Network
To understand the impact of the gating network, we ablate by removing it from MiDRE, reported in Table 4 as w/o Gate. In this case, we simply add the output from the internal reasoning expert and external reasoning expert. We notice that performance drops notably (Acc: $\downarrow 3.26\%$ in MMSD and $\downarrow 1.34\%$ in MMSD2.0). This decrease in performance indicates that the gating network plays a crucial role in controlling the interaction between the internal reasoning network (IR) and external reasoning module (ER). The adaptive gating ensures that the model relies on the most informative reasoning path to make a decision. In cases where it can’t rely on one single reasoning path, it happens to look at both reasoning paths to arrive at a decision, which is described in detail in Section 7. We also validate our design choice of using a bottleneck architecture for the gating mechanism by comparing it against a simpler linear gate, as reported in Table 4 (w Linear Gate). The bottleneck gate consistently performs better, suggesting that it facilitates more expressive interactions between the expert outputs.
@@ 6.3. Need of All Three Rationales
To validate the necessity of all three rationales (text, image and multimodal) in MiDRE, we conduct a systematic ablation study presented in Table 4 (rows 1 & 2). When only the multimodal rationale is used, the model’s performance drops significantly,(Acc: $\downarrow 3.68\%$ in MMSD and $\downarrow 2.12\%$ in MMSD2.0), indicating that relying solely on the multimodal rationale is insufficient. Similarly, when only the text and image rationale is used, the performance drops (Acc: $\downarrow 3.14\%$ in MMSD and $\downarrow 1.5\%$ in MMSD2.0), compared to the full model. In contrast, the full MiDRE model, which combines the image, text and multimodal rationales achieves the highest performance , emphasizing the necessity of views from both image and text modalities as well as the view from the interaction between them.
7. Qualitative Analysis
We qualitatively explore the following research questions: (1) How does external rationale help? (2) How does noisy external rationale affect model performance?
MiDRE MMSD MMSD2.0 Acc. P R F1 Acc. P R F1 Only Mul. Rationale 90.13 87.35 90.41 88.85 84.06 80.51 86.69 83.48 Only Txt+Img Rationale 90.67 89.21 88.37 88.78 84.68 82.11 82.35 82.23 w/o ER 89.09 85.82 89.11 87.43 82.27 74.51 89.39 81.28 w/o IR 91.06 89.62 88.45 89.03 84.06 80.03 83.89 81.92 w/o Gate 90.55 85.13 89.97 87.48 84.84 78.96 88.33 83.38 w Linear Gate 91.11 85.82 90.02 87.86 84.89 79.17 88.52 83.58 MiDRE 93.81 93.11 93.84 93.47 86.18 86.62 89.00 87.79

### v2

@@ 3.5. Model Training and Prediction
The output $H_{P}^{L}$ from the final encoder layer $L$ is fed to the LLM decoder, which generates the label from the label set $\{yes,no\}$. The LLM and the image and text projection layers are trained to minimize the cross-entropy loss between the ground truth label and the generated label. The CLIP visual and text encoders are only used as feature extractors and are frozen during training.
@@ 4.1. Datasets and Experimental Settings
We evaluate our proposed model MiDRE on MMSD2.0 (Qin et al., 2023) dataset. We also report the performance on MMSD(Cai et al., 2019), which is a biased dataset in Table 3 and discussed in §4.4. The dataset statistics are reported in Table 1. We use flan-t5-large [^1] https://huggingface.co/google/flan-t5-large as our backbone LLM. For generating rationales, we use the LVLM Qwen2-VL-7B-Instruct [^2] https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct, due to its better zero-shot performance on our task over other LVLMs as validated in Figure 4. For image and text feature extractors, we use clip-vit-large-patch14333https://huggingface.co/openai/clip-vit-large-patch14. We train our model for 20 epochs with Adam optimizer, maintaining a batch size of 12 and a learning rate of 5e-5. We pick the checkpoint with the best validation accuracy for testing. All experiments are carried out on 1xNvidia A100 80GB GPU.
@@
Negative 11174/10230 1451/1368 1450/1372
Figure 4. Zero-shot Performance of LVLMs on MMSD2.0 test samples.
Table 2. Performance comparison of different models on the MMSD2.0 dataset. We reimplement all models. indicates our method significantly outperforms baselines with $p<0.05$. Best results are in bold.
Model Acc. P R F1 Text-Only TextCNN (Kim, 2014) 71.61 64.62 75.22 69.52 BiLSTM (Graves and Schmidhuber, 2005) 72.48 68.02 68.08 68.05 SMSD (Xiong et al., 2019) 73.56 68.45 71.55 69.97 RoBERTa (Liu et al., 2019) 79.66 76.74 75.70 76.21 Image-Only ResNet (He et al., 2015) 65.50 61.17 54.39 57.58 ViT (Dosovitskiy et al., 2020) 72.02 65.26 74.83 69.72 Multimodal HFM (Cai et al., 2019) 70.57 64.84 69.05 66.88 Att-BERT (Pan et al., 2020) 80.03 76.28 77.82 77.04 CMGCN (Liang et al., 2022) 79.83 75.82 78.01 76.90 HKE (Liu et al., 2022b) 76.50 73.48 71.07 72.25 DynRT-Net (Tian et al., 2023) 71.40 71.80 72.17 71.34 Multi-view CLIP (Qin et al., 2023) 85.14 80.18 88.21 84.00 RAG-LLaVA (Tang et al., 2024) 85.11 80.02 87.34 83.51 MiDRE (ours) 86.18 86.62 89.00 87.79
@@ 4.2. Baselines
We consider three groups of baselines: image-modality, text-modality and multimodal methods. For image-modality, we follow (Cai et al., 2019) and use the representations from the pooling layer of ResNet (He et al., 2015) and embedding of the [CLS] token from ViT (Dosovitskiy et al., 2020). For text-modality methods, we consider TextCNN (Kim, 2014), BiLSTM (Graves and Schmidhuber, 2005), SMSD (Xiong et al., 2019) which uses a low rank bilinear pooling using a self-matching network, and RoBERTa(Liu et al., 2019). For the multimodal methods, we consider approaches designed for multimodal sarcasm detection as baselines, namely, HFM (Cai et al., 2019), D&R Net (Xu et al., 2020), Att-BERT (Pan et al., 2020), InCrossMGs (Liang et al., 2021), CMGCN (Liang et al., 2022), HKE (Liu et al., 2022a), DynRT-Net (Tian et al., 2023), Multi-view CLIP (Qin et al., 2023) and RAG-LLaVA (Tang et al., 2024). We do not show the results of InCrossMGs and D&R Net on MMSD2.0 due to the unavailability of their codes.
Table 3. Performance comparison of different models on the MMSD dataset. We reimplement all models. indicates our method significantly outperforms baselines with $p<0.05$. Best results are highlighted in bold.
Model Acc. P R F1 Text-Only TextCNN (Kim, 2014) 80.03 74.29 76.39 75.32 BiLSTM (Graves and Schmidhuber, 2005) 81.90 76.66 78.42 77.53 SMSD (Xiong et al., 2019) 80.90 76.46 75.18 75.82 RoBERTa (Liu et al., 2019) 93.97 90.39 94.59 92.45 Image-Only ResNet (He et al., 2015) 64.76 54.41 70.80 61.53 ViT (Dosovitskiy et al., 2020) 67.83 57.93 70.07 63.40 Multimodal HFM (Cai et al., 2019) 83.44 76.57 84.15 80.18 D&R Net (Xu et al., 2020) 84.02 77.97 83.42 80.60 Att-BERT (Pan et al., 2020) 86.05 80.87 85.08 82.92 InCrossMGs (Liang et al., 2021) 86.10 81.38 84.36 82.84 CMGCN (Liang et al., 2022) 86.54 - - 82.73 HKE (Liu et al., 2022b) 87.36 81.84 86.48 84.09 DynRT-Net (Tian et al., 2023) 93.59 93.06 93.60 93.31 Multi-view CLIP (Qin et al., 2023) 88.33 82.66 88.65 88.55 RAG-LLaVA (Qin et al., 2023) 89.97 89.26 89.58 89.42 MiDRE (ours) 95.81 93.11 93.84 93.47
@@ 4.3. Main Results
Following (Cai et al., 2019), we use Accuracy (Acc.), macro-average precision (P), macro-average recall (R), and macro-average F1 (F1) to report the performance of our model in Table 2. Our main observations are follows: (1) MiDRE outperforms all SOTA baselines in MMSD2.0: On MMSD2.0 dataset, MiDRE outperforms all baseline across unimodal and multimodal settings. Specifically it ourpeforms the best SOTA multimodal baseline Multi-view CLIP by good margins across metrics (Acc: $\uparrow 1.04\%$, P: $\uparrow 6.44\%$, R: $\uparrow 0.79\%$, F1: $\uparrow 3.79\%$ ). This proves the efficacy of using structured rationales and having dual reasoning paths to identify sarcasm.
(2) MiDRE outperforms all multimodal baselines that use external knowledge: We notice that our method significantly outperforms baselines like HFM (Cai et al., 2019) (Acc: $\uparrow 15.61\%$ in MMSD2.0) which uses image attributes generated using ResNet, CMGCN (Liang et al., 2022) (Acc: $\uparrow 6.35\%$ in MMSD2.0) which utilizes object-attribute pairs from a detectron module and HKE (Liu et al., 2022a) (Acc: $\uparrow 9.68\%$ in MMSD2.0) which incorporates image caption, image attributes and adjective-noun pairs from image. This highlights the importance of leveraging structured, verbose and sarcasm-oriented knowledge.
@@ 4.4. Results on Biased MMSD Dataset
MMSD is a text-biased dataset, as shown by Qin et al. (2023), due to strong textual cues like hashtags and emojis that predominantly occur in one class. Text-only models can exploit these cues without considering visual input, often outperforming multimodal models. For completeness, we evaluate MiDRE on MMSD. As shown in Table 3, the higher performance of RoBERTa over multimodal baselines highlights this bias. Nonetheless, MiDRE consistently outperforms all unimodal and multimodal models by significant margins, demonstrating robust reasoning ability.
Table 4. Performance comparison of MiDRE with LVLMs on MMSD2.0. Best scores are in bold.
Model Acc P R F1 MiDRE (Ours) 86.18 86.62 89.00 87.79 Qwen-2-VL-7B (Yang et al., 2024) 85.71 83.36 86.20 84.69 LLaMA-3.2-11B (Dubey et al., 2024) 84.64 80.20 85.26 82.56 LLaVA-1.6-7B (Liu et al., 2023) 85.38 77.48 93.15 84.54 InternVL-2.5-8B (Chen et al., 2024) 85.26 78.21 91.13 84.12
MiDRE MMSD MMSD2.0 Acc. P R F1 Acc. P R F1 Only Mul. Rationale 90.13 87.35 90.41 88.85 84.06 80.51 86.69 83.48 Only Txt+Img Rationale 90.67 89.21 88.37 88.78 84.68 82.11 82.35 82.23 w/o ERE 89.09 85.82 89.11 87.43 82.27 74.51 89.39 81.28 w/o IRE 91.06 89.62 88.45 89.03 84.06 80.03 83.89 81.92 w/o Gate 90.55 85.13 89.97 87.48 84.84 78.96 88.33 83.38 w Linear Gate 91.11 85.82 90.02 87.86 84.89 79.17 88.52 83.58 w Swap Enc (P $\leftrightarrow$ T) 91.62 89.23 88.75 88.99 82.13 81.62 80.20 80.90 MiDRE 95.81 93.11 93.84 93.47 86.18 86.62 89.00 87.79
Figure 5. Unique test samples correctly classified by MiDRE(w/o IRE) vs MiDRE(w/o ERE).
@@ 5. Comparison with LVLMs
We compare the performance of MiDRE with four SOTA large vision-language models (LVLMs) by LoRA finetuning them with rank 8, due to resource constraints. The results in Table 4 show that MiDRE consistently outperforms all LVLMs across most metrics. While LLaVA-1.6-7B records the highest recall (93.15%), its lower precision leads to a weaker F1 than MiDRE. This suggests that MiDRE not only identifies relevant instances effectively but also minimizes false positives, making it more robust. At the same time, it reveals how the capabilities of such smaller models can be significantly amplified by strategically leveraging contextual understanding from large vision-language models.
Figure 6. Examples demonstrating the importance of external rationales for sarcasm identification. ’Baseline’ indicates Multi-view CLIP (the best performing baseline for MMSD2.0), ERE(IRE) Gate Values represent the importance given to external (internal) expert module by the gating network in MiDRE.
@@ 6. Ablation
To further prove the efficacy of our approach, we answer the following research questions: (1) What is the impact of having only one expert network? (2) What advantage does the gating network provide? (3) Do we need rationales from all of image, text and multimodal views?
@@ 6.1. Single Expert Network
To better understand the contributions of IRE and ERE, we perform an ablation study on MiDRE by selectively removing each component (Table 5, rows 3 & 4). Removing the ERE module (w/o ERE) leads to a significant drop in accuracy (↓6.72% on MMSD, ↓3.91% on MMSD2.0), highlighting the importance of external contextual rationales that are not directly inferable from the input. Conversely, removing IRE (w/o IRE) also reduces performance (↓4.75% on MMSD, ↓2.12% on MMSD2.0), showing that ERE alone struggles to compensate for noisy rationales.
Interestingly, from Figure 5, we can observe that w/o ERE (only considering IRE) correctly classifies 111 unique samples and w/o IRE (only considering ERE) classifies 154 different ones, suggesting that both experts offer complementary and independently valuable capabilities. Moreover, the larger performance drop without ERE indicates its dominant role, but IRE remains essential for grounding predictions in the raw image-text content.
Refer to caption Refer to caption Text: # bringbackobama because who needs the constitution anyway Text: the beautiful los angeles river-why isn’t this a wonder of the world already? Label: Sarcastic Label: Sarcastic Type: External Type: Internal Refer to caption Refer to caption Text: radicalizing my biceps! Text: i ’m crying this is so cute Label: Non-sarcastic Label: Non-sarcastic Type: External Type: Internal
Figure 7. Representative examples. Top row: internal vs. external-knowledge sarcastic cases. Bottom row: internal vs. external-knowledge non-sarcastic cases.
Figure 8. Gating values distribution across 4 categories of samples. External means samples that heavily rely on external knowledge, while internal represent those classes of samples that do not require external context.
Figure 9. Examples demonstrating how noisy rationales can provide contextual cues. ’Baseline’ indicates Multi-view CLIP (the best performing baseline for MMSD2.0), ERE(IRE) Gate Values represent the importance given to external (internal) expert module by the gating network in MiDRE.
@@ 6.2. Impact of Gating Network
To understand the impact of the gating network, we ablate by removing it from MiDRE, reported in Table 5 as w/o Gate. In this case, we simply add the output from IRE and ERE. We notice that performance drops notably (Acc: $\downarrow 5.26\%$ in MMSD and $\downarrow 1.34\%$ in MMSD2.0). This decrease in performance indicates that the gating network plays a crucial role in controlling the interaction between the IRE and ERE. The adaptive gating ensures that the model relies on the most informative reasoning path to make a decision. In cases where it can’t rely on one single reasoning path, it happens to look at both reasoning paths to arrive at a decision, which is described in detail in Section 7. We also validate our design choice of using a bottleneck architecture for the gating mechanism by comparing it against a simpler linear gate, as reported in Table 5 (w Linear Gate). The bottleneck gate consistently performs better, suggesting that it facilitates more expressive interactions between the expert outputs.
@@ 6.3. Need of All Three Rationales
To validate the necessity of all three rationales (text, image and multimodal) in MiDRE, we conduct a systematic ablation study presented in Table 5 (rows 1 & 2). When only the multimodal rationale is used, the model’s performance drops significantly,(Acc: $\downarrow 5.68\%$ in MMSD and $\downarrow 2.12\%$ in MMSD2.0), indicating that relying solely on the multimodal rationale is insufficient. Similarly, when only the text and image rationale is used, the performance drops (Acc: $\downarrow 5.14\%$ in MMSD and $\downarrow 1.5\%$ in MMSD2.0), compared to the full model. In contrast, the full MiDRE model, which combines the image, text and multimodal rationales achieves the highest performance, emphasizing the necessity of knowledge from all three views.
7. Analysis
7.1. Gating Values Distribution
To assess whether the gating network prioritizes the ERE or IRE according to genuine reasoning needs rather than spurious cues, we perform a controlled experiment. We curated 100 test samples (50 sarcastic, 50 non-sarcastic). Within each class, we further identified 25 examples that require external knowledge (e.g., commonsense facts, cultural references, or world events) and 25 examples that can be resolved purely through internal image–text incongruity. The representative examples of each group is shown in Fig. 7. For each group, we recorded the gate’s softmax weights. We observe in Fig 8 that samples needing external knowledge showed consistently higher ERE weights across sarcasm labels, while internal-only cases favored IRE, suggesting that the gate aligns with reasoning needs rather than superficial cues.
7.2. Perturbation Analysis of Rationales
To assess whether MiDRE leverages the external rationales meaningfully rather than ignoring them or learning spurious correlations, we perform a controlled perturbation experiment. We select the same 50 samples (25 sarcastic and 25 non-sarcastic) used in §7.1, which require external knowledge. For each sample, we replace its original rationale with a rationale randomly drawn from another sample in this set (ensuring no self-substitution), while keeping the image and text unchanged. We then re-run the full pipeline and measure: (1) classification accuracy before and after perturbation, and (2) the gate’s value for ERE, before and after perturbation.
Table 6 shows the effect of perturbing external rationales on classification accuracy and the average ERE gating weight. Across all external-knowledge samples, accuracy drops by $\downarrow 4.7\%$, with a slightly larger decrease for sarcastic cases ($\downarrow 5.2\%$) than for non-sarcastic cases ($\downarrow 4.2\%$), reflecting greater reliance on external knowledge when sarcasm is present. The mean gate value for ERE also decreases across groups (e.g., $\downarrow 0.06$ for sarcastic samples), indicating that the gating network appropriately down-weights the ERE when its reasoning is unreliable. However, since these posts genuinely require external knowledge, this reduction in reliance leads to performance degradation, demonstrating that the gate’s behavior aligns with reasoning needs rather than spurious cues.
Table 6. Perturbation analysis of the external rationales. Negative $\Delta$ shown in red.
Group # Accuracy (%) ERE Gate Value
Orig Pert $\Delta$ Orig Pert
All ext.-knowledge 50 88.3 83.6 4.7-4.7 0.70 0.65
Sarcastic 25 89.3 84.1 5.2-5.2 0.74 0.68
Non-sarcastic 25 87.4 83.2 4.2-4.2 0.67 0.62